OpenStudy (anonymous):

Hi, I was having some trouble with natural logarithms. Does anybody know how to solve [3^(2x)+4(3^(x+1)) = 189] for x using natural logarithms? I know it's a monster, sorry. :P

OpenStudy (anonymous):

make substitution 3^x=y. It will be a quadratic.

OpenStudy (anonymous):

Sorry, but would you mind explaining that to me? So I write 3^(2X) + 4y = 189, and then what?

OpenStudy (anonymous):

No, the eqn will be y^2+4(3)y=189. Hope u can figure out why.:)

OpenStudy (anonymous):

Oh, I see! I think I can manage now. Thanks! :)

OpenStudy (anonymous):

4the second term is 12 multiplied by 3 to the x

OpenStudy (anonymous):

you let 3 raised to x be p, then you end up with a quadractic equation p squared plus tweleve p equals 189

OpenStudy (anonymous):

p\[p ^{2}+12p=189\]